# TreeDGS: Aerial Gaussian Splatting for Distant DBH Measurement

### v2

@@ 4.1 Experimental settings
We evaluate TreeDGS on the 10 circular field plots described in Section 2. Field DBH was measured with a diameter tape at breast height ($h_{\mathrm{BH}}=1.37$ m above ground, Section 2). We compare direct DBH measurement under the following reconstruction sources and fitting strategies: (i) UAV LiDAR: DBH estimation from the aerial LiDAR point cloud using (a) cylinder fitting [32] and (b) our slice-wise circle fitting (non-weighted and intensity-weighted variants). (ii) TreeDGS: DBH estimation from the point cloud sampled from the optimized 3D Gaussian field (Section 3.4), using the same cylinder and circle fitting variants, with opacity-based weighting available only for TreeDGS. To ensure a fair comparison, trunk/instance segmentation for both LiDAR and TreeDGS inputs is performed using the same ForestFormer3D checkpoint [48].
Cylinder fitting baseline.
@@
We follow DigiForests [32] for cylinder fitting. A cylinder is fit to each trunk point cloud using a RANSAC-initialized model. Each RANSAC round samples five points, estimates the cylinder center and axis from the covariance SVD, and keeps models whose axis is within $\pi/8$ of vertical and radius $\leq 0.5\,\mathrm{m}$. The RANSAC cylinder is then refined by least squares, updating the cylinder center, axis, and radius to minimize a robust (Geman–McClure) sum of squared point-to-cylinder distance residuals. Iterations stop when the parameter update norm falls below $\lVert\Delta x\rVert<10^{-5}$, where $\Delta x\in\mathbb{R}^{7}$ stacks the center shift, rotation/axis update, and radius change. If the least-squares refinement increases the inlier ratio, it replaces the RANSAC estimate; otherwise the RANSAC cylinder is retained. Finally, the diameter at breast height is reported as $\mathrm{DBH}=2r$ from the selected cylinder radius.
Figure 7: LiDAR intensity vs. TreeDGS opacity as a reliability signal at a same 5m slice height (Plot 2). (a) UAV LiDAR point cloud colored by normalized return intensity. (b) TreeDGS surface samples colored by learned opacity (0–1). Bottom: zoomed top-down views of two example stems ((a-1),(a-2) intensity; (b-1),(b-2) opacity). Compared to LiDAR intensity, the opacity values form a clearer high-confidence core around the stem cross-sections. This provides per-point confidence cue for DBH circle fitting. LiDAR has 281.4K points and TreeDGS has 66.6K points.
Figure 8: Qualitative comparison of point density and trunk completeness (Plot 1). (a) SfM points are sparse with limited stem coverage. (b) OpenMVS densification improves density but remains incomplete on trunks under canopy occlusion. (c) TreeDGS extracts a dense point set from the optimized Gaussians via opacity-guided sampling and RaDe-GS depth-aware cumulative-opacity tagging, yielding more continuous trunk support near breast height. (Max 100 candidate draws per splat; retained points vary with opacity.) (d) UAV LiDAR provides strong structure but can still be stem-sparse/contaminated depending on scan geometry and occlusion. Colors show height above ground.
Circle fitting variants.
@@
For the proposed circle fitting, we evaluate two reliability modes: non-weighted (nw), which treats all points equally, and weighted (w), which uses a per-point reliability score as a weight (Eq. (10)) in both hypothesis sampling and inlier scoring (Eq. (14)). For TreeDGS, this reliability is the multi-view opacity consistency estimated by our renderer. For LiDAR, we additionally report an intensity-weighted analogue that uses normalized return intensity as a proxy weight. In our UAV data, this does not improve DBH accuracy because raw airborne LiDAR intensity is often speckled and unstable without careful radiometric correction and calibration (e.g., sensitivity to range, incidence angle, and scanner settings) [12, 21, 19, 46, 49]. Fig. 7 visualizes this gap: opacity yields a cleaner, geometry-aligned confidence core around stem cross-sections compared to noisy intensity. In addition, we omitted SfM/MVS benchmarking because breast-height trunk points were insufficient for stable segmentation and fitting in this dataset (See Fig. 8).
Across all circle-fitting variants, we use slice thickness $H=1.0$ m and slice spacing $\Delta z=0.1$ m, discarding slices with fewer than $5$ points. For each slice, we run solid-circle RANSAC (Eq. (14)) with $K=2000$ hypotheses, a minimum inlier fraction $\rho_{\min}=0.1$, and radius bounds $r_{\min}=0.02$ m and $r_{\max}=1.0$ m. We set the radius regularization exponent to $p=0.6$ via held-out validation on a 10% subset of the matched data and keep it fixed for all experiments, reporting results only on disjoint test data. The taper model in Eq. (15) uses RANSAC with residual threshold $\epsilon=2$ cm (max trials $T=1000$, min samples $3$), requiring at least $10$ inlier slices and enforcing $\beta_{1}<0$. Finally, DBH is reported as $\mathrm{DBH}=\hat{d}(h_{\mathrm{BH}})$ (Eq. (16)) evaluated at $h_{\mathrm{BH}}=1.37$ m.
@@
We report four error measures between estimated and field DBH: root mean squared error (RMSE), relative RMSE (RRMSE, normalized by the mean field DBH), mean absolute error (MAE), and mean error (ME, bias). All errors are in cm, lower is better for RMSE/RRMSE/MAE, and ME closer to 0 indicates lower systematic bias.
Figure 9: Estimated vs. field DBH by plot (all methods). Each panel shows per-tree predictions versus field tape measurements for one plot; the dashed line is the 1:1 reference. Across most plots, TreeDGS with opacity-weighted circle fitting clusters closer to the 1:1 line than unweighted fits and the TreeDGS+cylinder baseline, indicating reduced variance and substantially lower bias.
Figure 10: Mean-only vs. proposed point sampling from Gaussian Splats. (a) Exporting only Gaussian means [27] (one point per splat) yields an extremely sparse, fragmented cloud that is insufficient for downstream tasks such as 3D segmentation (ForestFormer3D [48]). (b) Our opacity-guided surface sampling densifies geometry and preserves stem surfaces and canopy structure.
@@ 4.2 Results
Table 2 summarizes per-plot performance and includes a pooled All row. TreeDGS with opacity-weighted circle fitting achieves the best overall accuracy (pooled RMSE =4.79=4.79 cm; MAE =3.70=3.70 cm) with negligible bias (ME =0.38=-0.38 cm), while maintaining a high success rate (189/190). Compared with the UAV LiDAR cylinder-fitting baseline, TreeDGS reduces pooled RMSE by approximately 39% while using only RGB imagery.
@@
Ablations also show that the fitting model must match the reconstructed geometry. Applying a classical cylinder fit [32] to points sampled from TreeDGS performs poorly (13.29 cm RMSE; ME =6.46=-6.46 cm), consistent with our discussion in Section 3.6: TreeDGS sampling produces slice-wise cross-sections rather than a thin ring of surface points. Similarly, unweighted circle fitting is substantially less accurate (9.61 cm RMSE), showing that down-weighting low-confidence samples using multi-view opacity is a better approach.
TreeDGS (Cir. w. fit.) attains the lowest RMSE in most plots (Plots 1, 3–6, 8–9) and is within 0.34 cm RMSE of the LiDAR cylinder-fitting baseline in Plot 2 (Table 2). The largest gap appears in Plot 3: LiDAR (cyl. fit.) reaches 17.22 cm RMSE, whereas TreeDGS (Cir. w. fit.) remains at 5.93 cm, indicating greater robustness of alpha/opacity weighted in circle fitting.
Fig. 9 visualizes per-tree DBH errors. TreeDGS (cyl. fit.) exhibits a consistent underestimation trend (points below the 1:1 line), matching the negative pooled bias in Table 2. Unweighted circle fitting (TreeDGS (cir. nw. fit.)) shows greater scatter and occasional large overestimation, consistent with low-confidence samples inflating the fitted disk. In contrast, the opacity-weighted fit (TreeDGS (cir. w. fit.)) is more tightly clustered around the 1:1 line, indicating that multi-view rendering consistency provides a useful reliability cue for suppressing floating samples and occlusion-induced artifacts.
@@
We additionally report a simple baseline that exports one point per Gaussian by using the splat centers $\{\boldsymbol{\mu}_{i}\}$, similar to trunk-focused usage in prior 3DGS tree pipelines [27]. In our stand-off UAV forest scenes this produces an extremely sparse and surface-incomplete point set (Fig. 10), which breaks downstream instance segmentation with ForestFormer3D [48] and prevents reliable DBH fitting. These failures highlight that surface-densified sampling is necessary in the stand-off UAV setting.
5 Conclusion
We introduced TreeDGS: a UAV RGB-only pipeline that repurposes 3D Gaussian Splatting as an opacity-weighted density sampling method for trunk measurement, rather than only a view-synthesis renderer. Building on SfM/MVS initialization, TreeDGS optimizes a Gaussian field and then converts it into a measurement-ready stem point set via depth-aware cumulative-opacity sampling and multi-view reliability weighting, enabling robust circle fitting at breast height. This approach bridges a key gap left by prior UAV LiDAR and UAV-SfM DBH pipelines that operate on inherently sparse or view-limited point clouds [35, 34], and demonstrates that Gaussian splats can support accurate trunk-level metrics from commodity aerial imagery [22].

### v3

@@ 4.1 Experimental Settings
We evaluate TreeDGS on the 10 circular field plots described in Section 2. Field DBH was measured with a diameter tape at breast height ($h_{\mathrm{BH}}=1.37$ m above ground, Section 2). We compare direct DBH measurement under the following reconstruction sources and fitting strategies: (i) UAV LiDAR: DBH estimation from the aerial LiDAR point cloud using (a) cylinder fitting [33] and (b) our slice-wise circle fitting (non-weighted and intensity-weighted variants). (ii) TreeDGS: DBH estimation from the point cloud sampled from the optimized 3D Gaussian field (Section 3.4), using the same cylinder and circle fitting variants, with opacity-based weighting available only for TreeDGS. To ensure a fair comparison, trunk/instance segmentation for both LiDAR and TreeDGS inputs is performed using the same ForestFormer3D checkpoint [50]. If ForestFormer3D misses a matched tree instance (e.g., due to strong bending/occlusion) or if the downstream fitting cannot return a DBH estimate, we count that tree as a failure when computing the Success Rate (SR) in Table 2.
@@
Cylinder fitting baseline. We follow DigiForests [33] for cylinder fitting. A cylinder is fit to each trunk point cloud using a RANSAC-initialized model. Each RANSAC round samples five points, estimates the cylinder center and axis from the covariance SVD, and keeps models whose axis is within $\pi/8$ of vertical and radius $\leq 0.5\,\mathrm{m}$. The RANSAC cylinder is then refined by least squares, updating the cylinder center, axis, and radius to minimize a robust (Geman–McClure) sum of squared point-to-cylinder distance residuals. Iterations stop when the parameter update norm falls below $\lVert\Delta x\rVert<10^{-5}$, where $\Delta x\in\mathbb{R}^{7}$ stacks the center shift, rotation/axis update, and radius change. If the least-squares refinement increases the inlier ratio, it replaces the RANSAC estimate; otherwise, the RANSAC cylinder is retained. Finally, the diameter at breast height is reported as $\mathrm{DBH}=2r$ from the selected cylinder radius.
@@
Circle fitting variants. For the proposed circle fitting, we evaluate the following two reliability modes: non-weighted (nw), which treats all points equally, and weighted (w), which uses a per-point reliability score as a weight (Equation (10)) in both hypothesis sampling and inlier scoring (Equation (14)). For TreeDGS, this reliability is the multi-view opacity consistency estimated by our renderer. For LiDAR, we additionally report an intensity-weighted analogue that uses normalized return intensity as a proxy weight. In our UAV data, intensity-weighting helps reduce the strong positive bias of non-weighted solid-circle fitting (Table 2), but it still underperforms LiDAR cylinder fitting and has lower SR, consistent with raw airborne LiDAR intensity being speckled and unstable without careful radiometric correction and calibration (e.g., sensitivity to range, incidence angle, and scanner settings) [12, 22, 20, 48, 51].
Figure 8 visualizes this gap: opacity yields a cleaner, geometry-aligned confidence core around stem cross-sections compared to noisy intensity. In addition, we omitted SfM/MVS benchmarking because breast-height trunk points were insufficient for stable segmentation and fitting in this dataset (Figure 9).
Across all circle-fitting variants, we use slice thickness $H=1.0$ m and slice spacing $\Delta z=0.1$ m, discarding slices with fewer than $5$ points. For each slice, we run solid-circle RANSAC (Equation (14)) with $K=2000$ hypotheses, a minimum inlier fraction $\rho_{\min}=0.1$, and radius bounds $r_{\min}=0.02$ m and $r_{\max}=1.0$ m. We select the radius regularization exponent $p$ in Equation (14) via held-out validation on a random selection of 10% of the matched stems and then keep it fixed for all test experiments. We use the same validation/test split IDs for all compared methods for fair comparison. Concretely, from the 210 matched trees in Table 1, we randomly reserve 21 trees (10%) for validation and report all numbers in Table 2 on the remaining disjoint test set (N=189). This is why the per-plot denominators in Table 2 differ from Table 1. We obtain $p=0.6$ for TreeDGS opacity weighting and $p=0.85$ for the LiDAR intensity-weighted analogue (Figure 10).
The taper model in Equation (15) uses RANSAC with residual threshold $\epsilon=2$ cm (max trials $T=1000$, min samples $3$), requiring at least $10$ inlier slices and enforcing $\beta_{1}<0$. Finally, DBH is reported as $\mathrm{DBH}=\hat{d}(h_{\mathrm{BH}})$ (Equation (16)) evaluated at $h_{\mathrm{BH}}=1.37$ m.
Figure 8: LiDAR intensity vs. TreeDGS opacity as a reliability signal at a same 5m slice height (Plot 2). (a) UAV LiDAR point cloud colored by normalized return intensity. (b) TreeDGS surface samples colored by opacity (0–1). Bottom: zoomed top-down views of two example stems ((a-1,a-2) intensity; (b-1,b-2) opacity). Compared to LiDAR intensity, the opacity values form a clearer high-confidence core around the stem cross-sections. This provides per-point confidence cue for DBH circle fitting. LiDAR has 281.4K points and TreeDGS has 66.6K points.
Figure 9: Qualitative comparison of point density and trunk completeness (Plot 1). (a) SfM points are sparse with limited stem coverage. (b) OpenMVS densification improves density but remains incomplete on trunks under canopy occlusion. (c) TreeDGS extracts a dense point set from the optimized Gaussians via opacity-guided sampling ($M=100$) and RaDe-GS depth-aware cumulative-opacity tagging, yielding more continuous trunk support near breast height. (d) UAV LiDAR provides strong structure but can still be stem-sparse/contaminated depending on scan geometry and occlusion. Colors indicate height above ground.
Figure 10: Held-out validation for selecting the radius penalty exponent $p$ in Equation (14). We sweep $p$ for weighted solid-circle RANSAC and measure RMSE/MAE on the held-out validation split. (a) TreeDGS opacity-weighted circle fitting selects $p=0.6$. (b) LiDAR intensity-weighted circle fitting selects $p=0.85$. We keep these values fixed when reporting all test results in Table 2.
@@
We report four standard error measures between estimated and field DBH: root mean squared error (RMSE), relative RMSE (RRMSE, normalized by the mean field DBH), mean absolute error (MAE), and mean error (ME, bias). All errors are reported in centimeters, and lower is better for RMSE/RRMSE/MAE while ME closer to 0 indicates lower systematic bias. We additionally report the Success Rate (SR) as the fraction of trees in the test split for which a method returns a valid DBH estimate (Table 2). RMSE/RRMSE/MAE/ME are computed over the successful subset, while SR captures failures (e.g., missing instances or insufficient inlier slices for the taper RANSAC).
Table 2: Per-plot DBH accuracy and robustness (All–5). RMSE/MAE/ME are in cm, RRMSE is in %, and SR is success rate. Abbrev.: L/T = LiDAR/TreeDGS; Cyl = cylinder; CN/CW = circle non-weighted/weighted.
Plot Meth. RMSE RRMSE MAE ME SR
All L-Cyl 7.66 28.09 5.23 1.50 189/189
L-CN 16.61 60.96 14.81 14.67 187/189
L-CW 10.29 36.97 8.79 1.02 176/189
T-Cyl 13.29 48.69 10.25 -6.46 189/189
T-CN 9.61 35.22 6.97 5.79 189/189
T-CW 4.79 17.54 3.70 -0.38 189/189
1 L-Cyl 8.16 26.14 5.71 -0.04 19/19
L-CN 10.12 32.42 8.57 7.57 19/19
L-CW 9.94 32.04 8.04 2.01 18/19
T-Cyl 11.69 37.47 10.99 -10.99 19/19
T-CN 7.58 24.29 5.91 2.63 19/19
T-CW 4.94 15.83 3.72 -2.04 19/19
2 L-Cyl 5.36 18.84 4.34 0.38 22/22
L-CN 13.27 46.66 11.92 11.92 22/22
L-CW 8.60 30.25 7.12 -0.59 22/22
T-Cyl 13.43 47.21 9.84 -5.49 22/22
T-CN 8.43 29.65 7.26 7.09 22/22
T-CW 5.07 17.82 3.18 1.88 22/22
3 L-Cyl 12.92 49.55 7.75 4.47 20/20
L-CN 20.89 80.11 18.69 18.69 20/20
L-CW 11.85 45.44 10.32 3.55 20/20
T-Cyl 10.49 40.23 8.77 -7.35 20/20
T-CN 16.39 62.83 9.47 8.15 20/20
T-CW 5.93 22.73 4.35 -0.10 20/20
4 L-Cyl 6.63 27.06 5.81 2.94 16/16
L-CN 19.34 78.87 17.95 17.95 16/16
L-CW 9.59 35.77 7.78 -1.41 12/16
T-Cyl 25.24 102.93 16.57 3.84 16/16
T-CN 11.46 46.73 7.59 4.95 16/16
T-CW 5.74 23.40 4.53 -0.62 16/16
5 L-Cyl 6.20 21.87 4.27 2.37 21/21
L-CN 17.30 60.96 16.60 16.60 20/21
L-CW 11.61 40.92 10.99 3.26 20/21
T-Cyl 12.39 43.69 11.01 -7.98 21/21
T-CN 5.84 20.60 4.64 3.57 21/21
T-CW 4.38 15.46 4.01 -2.10 21/21
Table 3: Per-plot DBH accuracy and robustness (cont., plots 6–10).
Plot Meth. RMSE RRMSE MAE ME SR
6 L-Cyl 4.96 18.41 3.27 1.27 22/22
L-CN 14.03 52.15 12.86 12.86 22/22
L-CW 9.45 35.12 7.60 0.26 22/22
T-Cyl 8.64 32.10 7.91 -7.45 22/22
T-CN 8.91 33.10 8.23 8.23 22/22
T-CW 3.15 11.72 2.48 0.91 22/22
7 L-Cyl 9.16 44.08 7.23 1.92 17/17
L-CN 15.79 75.98 14.86 14.55 17/17
L-CW 7.79 33.96 6.31 -2.32 13/17
T-Cyl 13.66 65.77 8.15 -0.28 17/17
T-CN 12.01 57.81 10.57 9.68 17/17
T-CW 5.91 28.44 5.03 3.27 17/17
8 L-Cyl 9.65 33.11 7.80 2.61 16/16
L-CN 20.61 70.70 18.21 18.21 16/16
L-CW 11.07 36.67 9.77 1.11 15/16
T-Cyl 11.83 40.59 10.97 -7.64 16/16
T-CN 5.42 18.59 3.96 3.83 16/16
T-CW 3.68 12.63 3.14 -1.69 16/16
9 L-Cyl 6.18 23.13 4.20 -0.38 16/16
L-CN 16.27 62.05 15.09 15.09 15/16
L-CW 11.08 42.26 10.33 1.92 15/16
T-Cyl 8.73 32.66 7.73 -7.73 16/16
T-CN 8.50 31.81 6.85 5.17 16/16
T-CW 3.93 14.72 3.08 -1.26 16/16
10 L-Cyl 3.79 12.84 3.00 -0.30 20/20
L-CN 16.95 57.41 14.93 14.80 20/20
L-CW 10.73 35.95 9.23 0.84 19/20
T-Cyl 12.07 40.88 11.40 -11.15 20/20
T-CN 6.28 21.27 5.15 4.10 20/20
T-CW 4.42 14.95 3.73 -2.35 20/20
@@ 4.2 Results
Table 2 summarizes per-plot performance and includes a pooled ‘All’ row. All results in Table 2 are reported on the disjoint test split (N = 189) after reserving 21 of the 210 matched trees (10%) for validation/hyperparameter selection (see Experimental settings). TreeDGS with opacity-weighted circle fitting achieves the best overall accuracy (pooled RMSE =4.79=4.79 cm; MAE =3.70=3.70 cm) with negligible bias (ME =0.38=-0.38 cm), while maintaining a high success rate (189/189). Compared with the UAV LiDAR cylinder-fitting baseline (pooled RMSE =7.66=7.66 cm), TreeDGS reduces pooled RMSE by $\approx 37\%$ while using only RGB imagery.
@@
Ablations also show that the fitting model must match the reconstructed geometry. Applying a classical cylinder fit [33] to points sampled from TreeDGS performs poorly (13.29 cm RMSE; ME =6.46=-6.46 cm), consistent with our discussion in Section 3.6: TreeDGS sampling produces slice-wise cross-sections rather than a thin ring of surface points. Similarly, unweighted circle fitting is substantially less accurate (9.61 cm RMSE), showing that down-weighting low-confidence samples using multi-view opacity is a better approach. On LiDAR, the same trend holds: non-weighted solid-circle fitting strongly overestimates (ME =14.67=14.67 cm), while intensity-weighted fitting with a radius penalty substantially reduces this bias (ME =1.02=1.02 cm), though it still underperforms LiDAR cylinder fitting.
TreeDGS (Cir. w. fit.) attains the lowest RMSE in nine of ten plots (all except Plot 10) and achieves the best pooled performance (All row). In Plot 2 it slightly outperforms the LiDAR cylinder baseline (5.07 vs. 5.36 cm RMSE). The largest gap appears in Plot 3, where LiDAR (cyl. fit.) reaches 12.92 cm RMSE while TreeDGS (Cir. w. fit.) remains at 5.93 cm. In Plot 10, LiDAR (cyl. fit.) remains the best (3.79 cm RMSE).
Although a solid-disk inlier model can in principle encourage larger radii when volumetric interior points dominate, in our data the dominant failure mode for non-weighted fitting is positive bias driven by sparse exterior outliers (foliage/floaters) inflating the disk radius. The weighted objective in Equation (14) counteracts this through (i) reliability weights (opacity/intensity) and (ii) the explicit radius penalty $r^{-p}$, which requires proportionally more weighted inlier support to justify a larger radius. After optimizing $p$ on held-out validation (Figure 10), the estimator transitions from the positive-bias regime to near-unbiased behavior; the small pooled negative ME of 0.38-0.38 cm for TreeDGS (Cir. w. fit.) indicates mild shrinkage when only a partial high-confidence arc of the trunk boundary is visible and is within typical measurement/registration noise.
Success Rate (SR) in Table 2 explicitly accounts for cases where a method cannot return a DBH estimate. For example, LiDAR intensity-weighted circle fitting fails on 13 test trees (SR 176/189) because the taper RANSAC requires at least 10 inlier slices; intensity weighting can amplify slice-to-slice diameter variability and reduce the number of inlier slices below this threshold. In contrast, the non-weighted LiDAR circle fit fails on only two low-density trees (SR 187/189), suggesting that most additional failures are due to sensitivity to noisy intensity rather than fundamental point scarcity.
Beyond reporting pooled RMSE/MAE, we add three diagnostics to characterize robustness and failure modes. First, Figure 11 plots pooled per-tree signed error distributions (estimate-field). The opacity-weighted solid-circle variant yields the tightest distribution centered near zero, consistent with its lowest pooled RMSE/MAE and near-zero ME in Table 2. In contrast, the non-weighted solid-circle fits show a pronounced positive shift and heavier right tails (radius inflation), while cylinder fitting on TreeDGS samples shifts negative (systematic underestimation). Second, Figure 12 reports signed error versus ground-truth DBH. Size-dependent bias is clearly visible for several baselines: TreeDGS (Cyl. fit.) increasingly underestimates as DBH grows (large negative mean error for the largest stems), and LiDAR (Cir. nw. fit.) maintains strong positive bias across sizes. TreeDGS (Cir. w. fit.) shows substantially reduced size dependence: its binned mean error stays within single-digit centimeters and close to zero for the mid-range DBH values that dominate our plots ($\sim$18–38 cm), while becoming more negative in the largest-DBH bin (which contains relatively few trees). Third, to quantify the view-scarcity/occlusion bottleneck, Figure 13 relates the absolute DBH error (TreeDGS (Cir. w. fit.)) to simple breast-height (BH) observability metrics computed from the calibrated RGB views: the number of unoccluded BH views, the mean projected trunk width in pixels, and their product. On the 180/189 test trees with at least one unoccluded BH view, we observe weak but consistently negative correlations, indicating that higher BH visibility is associated with lower error. Importantly, the largest outliers (>>10 cm) occur predominantly in the low-visibility regime (few unoccluded views and/or small projected trunk width), supporting our discussion that even with opacity reliability weighting, accurate DBH requires at least a few informative views around breast height. Figure 14 shows an example of the BH-band view selection used for this analysis.
Figure 11: Aggregated DBH error distributions. Kernel density estimates of signed per-tree DBH error (estimate-field, cm) pooled across all plots for each method. The vertical grey dashed line marks zero error (perfect agreement with field DBH).
Figure 12: DBH error versus DBH size (aggregated across plots). Signed per-tree error (estimate-field, cm) versus ground-truth DBH on the test split, aggregated across all plots. Semi-transparent dots represent individual trees. For each method, thick solid segments show a robust binned error trend. Dashed segments are only used to bridge DBH ranges with no observations, linking the nearest available bins without implying data exist inside the gap.
[…]
Figure 14: Example BH-band visibility classification used in Figure 13. We project a small trunk band around $h_{\mathrm{BH}}$ into each calibrated RGB view and classify the crop (red box) as unoccluded or occluded. (Top): views counted as unoccluded where bark is visible within the BH band. (Middle): views rejected as occluded because foliage covers the BH band. (Bottom): corresponding top-full-frame context, illustrating how small the BH band is relative to the canopy in stand-off UAV imagery.
@@
We additionally report a simple baseline that exports one point per Gaussian by using the splat centers $\{\boldsymbol{\mu}_{i}\}$, similar to trunk-focused usage in prior 3DGS tree pipelines [28]. In our stand-off UAV forest scenes, this produces an extremely sparse and surface-incomplete point set (Figure 15), which breaks downstream instance segmentation with ForestFormer3D [50] and prevents reliable DBH fitting. These failures highlight that surface-densified sampling is necessary in the stand-off UAV setting.
Figure 15: Mean-only vs. proposed point sampling from Gaussian Splats. (a) Exporting only Gaussian means (one point per splat) yields an extremely sparse, fragmented cloud that is insufficient for downstream tasks such as 3D segmentation. (b) Our opacity-guided surface sampling densifies geometry and preserves stem surfaces and canopy structure. Point colors encode height above ground (blue low to red/orange high).
4.3 Ablation Study
4.3.1 Sensitivity Analysis of Fitting Parameters
In this section, we analyze the sensitivity of the DBH estimation stage to the main design parameters that govern slice construction (Equation (13)) and robust aggregation across height (Equation (15)). We vary one parameter at a time, slice thickness $H$, minimum points per slice $n_{\min}$, taper RANSAC residual threshold $\epsilon$, taper slope bound $\kappa$, and whether solid-circle fitting uses opacity reliability weights ($w_{k}=\bar{\alpha}_{k}$) or uniform weights ($w_{k}\equiv 1$) while keeping all other settings fixed ($\Delta z=0.1$ m, $K=2000$, $\rho_{\min}=0.1$, $r\in[0.02,1.0]$ m, and the tuned radius exponent $p$ from Figure 10).
Table 4 shows that DBH accuracy is stable over a broad range of values. Thin slabs increase error due to sparse support, whereas overly thick slabs can mix residual clutter; $H=1.0$ m performs best in our data and nearby choices remain competitive. Results are largely insensitive once $n_{\min}\geq 5$ and $\epsilon\geq 2$ cm. Imposing a moderate taper-slope bound improves robustness, and disabling this bound increases both RMSE and positive bias. Finally, using opacity as a continuous reliability weight yields a large improvement compared to uniform weighting, confirming the importance of the opacity signal.
Table 4: Ablation on TreeDGS on the held-out test split (N=189). Default values are shown in bold. SR is 189/189 for all rows, so only MAE, RMSE, and ME (cm) are shown.
Param. Val. MAE RMSE ME
Slice $H$ (m) 0.2 5.19 6.62 -3.85
0.5 3.86 5.23 -1.57
1.0 3.70 4.79 -0.38
1.5 3.89 6.42 0.41
2.0 4.03 5.94 0.88
$n_{\min}$ 3 3.94 6.35 -0.06
5 3.70 4.79 -0.38
10 3.66 4.82 -0.36
20 3.65 4.84 -0.31
$\epsilon$ (cm) 1.0 4.15 6.63 -0.25
2.0 3.70 4.79 -0.38
3.0 3.53 4.77 -0.41
5.0 3.53 4.73 -0.27
$\kappa$ (cm/m) 0.1 3.78 4.94 -0.62
0.2 3.71 4.87 -0.43
0.3 3.70 4.79 -0.38
0.5 3.68 4.86 -0.10
unb. 5.01 7.93 2.26
$w_{k}$ opacity 3.70 4.79 -0.38
uniform 6.97 9.61 5.79
4.3.2 ForestFormer3D Segmentation Quality
Because our field dataset does not provide the point-wise trunk/instance ground truth, we quantify segmentation performance on a small but representative subset (Plot 2) by creating a hand-corrected reference annotation. Starting from the ForestFormer3D predictions, a single annotator manually fixed trunk false negatives/positives, ground mislabels, and instance-ID assignment errors for all matched trees in Plot 2 ($\approx$14 labeling hours; Figure 16). Table 5 summarizes trunk semantic accuracy (F1 =0.812=0.812, IoU =0.683=0.683).
To quantify the impact on DBH, we reran the fitting pipeline on Plot 2 using the hand-corrected labels as an oracle segmentation. This reduces TreeDGS (Cir. w. fit.) RMSE from 5.07 cm to 4.21 cm (MAE from 3.18 cm to 2.68 cm) on the same 22 matched trees (Table 6). This indicates that segmentation errors contribute to DBH error, but a substantial portion of the remaining error is due to reconstruction/visibility limitations rather than segmentation alone.
Figure 16: ForestFormer3D segmentation on Plot 2 and our hand-corrected reference. (a) Off-the-shelf ForestFormer3D trunk/ground predictions and instance IDs (trunk points in red). (b) Hand-corrected labels used for evaluation, obtained by editing the predictions to fix trunk false negatives/positives, ground mislabels, and instance-ID assignment errors. Creating this reference annotation for Plot 2 required $\approx$ 14 hours by a single annotator. Yellow boxes highlight representative regions where manual edits corrected misclassifications/instance IDs; trunk points are red, while non-trunk points are blue/gray and ground points are green.
Table 5: ForestFormer3D segmentation accuracy on Plot 2. Hand-corrected reference; trunk semantic segmentation over 2,396,986 points.
Metric Value Notes
Precision 0.760 TP=107,746; FP=34,038
Recall 0.871 FN=15,974
F1 0.812
IoU 0.683
Accuracy 0.979 TN=2,239,228
Table 6: Effect of segmentation quality on DBH accuracy (Plot 2, 22 trees). TreeDGS (Cir. w. fit.) with ForestFormer3D predictions versus hand-corrected oracle labels. MAE/RMSE/ME are in cm and RRMSE is in %.
Segmentation SR MAE RMSE RRMSE ME
ForestFormer3D 22/22 3.18 5.07 17.82 1.88
Oracle 22/22 2.68 4.21 14.81 1.49
4.3.3 Runtime Analysis
[…]
Table 7: Runtime breakdown per plot (Plot 2, 365 images). Feature matching and Gaussian optimization dominate runtime; segmentation and DBH fitting are comparatively fast.
Stage Time Resource
Feature matching (TopicFM) 50 min GPU
SfM + bundle adjustment (GLOMAP) 3 min CPU
OpenMVS densification 5 min CPU
Gaussian optimization (RaDe-GS) 20 min GPU
ForestFormer3D segmentation 5 min GPU
Slice fitting + taper aggregation 10 s CPU
Total $\approx$83 min
Since plots are independent, the pipeline is trivially parallelizable across plots (one plot per GPU).
5 Discussion
Overall, our results suggest that TreeDGS can convert pixel-limited UAV RGB imagery into a trunk-centric 3D representation that is sufficiently coherent for plot-scale DBH estimation, narrowing the performance gap to a UAV LiDAR baseline while using only low-cost imagery. Compared to conventional SfM/MVS densification, Gaussian optimization combined with opacity-guided sampling produces denser and more spatially consistent support near breast height, which directly benefits downstream cross-sectional fitting. The added breast-height observability analysis further indicates that DBH errors increase primarily in the low-visibility regime, reinforcing that view scarcity and occlusion are key bottlenecks in aerial DBH measurement. In addition, our Plot 2 hand-correction study highlights that segmentation quality contributes measurably to DBH accuracy, but this does not fully account for residual error, implying that reconstruction completeness and boundary visibility remain limiting factors even with improved labels. Finally, the runtime breakdown shows that the full pipeline is practically feasible at the plot level, and that the proposed slice fitting stage is computationally negligible compared to reconstruction and learning-based components. These findings collectively support TreeDGS as a practical RGB-only alternative (or complement) for rapid forest inventory under appropriate visibility conditions. However, several limitations remain.
Limitations.